OpenStudy (anonymous):

The diameter of a circle is 2.6 centimeters. What is the circle's area? Use 3.14 for π and round your answer to the nearest hundredth.

OpenStudy (anonymous):

5.31 sqare centimeters

OpenStudy (anonymous):

Thanks Arnab

OpenStudy (anonymous):

this is just like the last one. A = pi(r^2) r = diameter /2 = 2.6 / 2 = 1.3 A = 3.14 (1.3^2) A = 5.31 cm^2 (note two numbers after the decimal to round to the nearest hundredth)